# Unleashing Cheapfakes through Trojan Plugins of Large Language Models

### 01 Dec 2023 v1

@@ 5.2 Backdoor Evaluation
Takeaway 5: Our Trojan adapter exhibits little malicious behavior on clean data and has little influence for the LLM-judged response quality.
5.3 Defense Evaluation
Previous defenses [55, 56] primarily focus on classification LMs, which cannot be directly applied [57, 58]. Therefore, we propose two potential mitigations. The first mitigation aims to detect a Trojan adapter by examining the rank values in the adapters. The intuition is that Trojan adapters trained to generate predefined target output can contain an abnormally distributed rank value in weight matrix. The second mitigation aims to remove potential backdoors or poisoning attacks by continuously fine-tuning the adapters on clean data.
Figure 10: Distribution of the maximum singular values of the query and value matrices in LoRA. The round points present singular pair for benign adapters. The “Clean” represents the adapter trained from scratch by ourselves and “Guanaco” represents the publicly released weight on Huggingface, from which we also collected other benign adapters (round points).
@@ 5.3 Defense Evaluation
Detection. As LoRA manipulates a foundational model with a low rank matrix, the singular value of the adapter’s matrix can reveal whether the adapter is maliciously trained. In addition to our trained clean adapter, we manually collected LoRAs from Huggingface and compare them along with all of our attacked adapters. The common modules of LoRAs are on the query and value matrix. We therefore compute, through SVD decomposition, the singular value pair $(q_{s},v_{s})$, where $q_{s}$ (resp., $v_{s}$) is the highest singular value of the query (resp., value) matrix in LoRA. Figure 10 visualizes the singular value pairs of shallow, medium and deep layers in clean and the attacked adapters. Our Trojan adapters are closely distributed while the clean adapters are positioned throughout the diagonal. However, this does not guarantee that our attack can be detected. In effect, the clean adapters are trained using a different algorithm, hyperparamters and training data, so their highest singular value are (almost) uniformly distributed on the diagonal. Meanwhile, the LoRAs trained with same setting as our Trojan adapters (pinpointed by red arrows) have singular pairs closely located.
@@
Removal. To remove a potential backdoor or poison in a compromised model, the defender can continuously fine-tune the downloaded adapter with clean data. Ideally, the defender fine-tunes the adapter on data of the same distribution in order to preserve its original performance. Therefore, we fine-tune our attacked adapters on clean OASST1 data. Figure 11 plots the KMR and RougeL scores for continuously fine-tuning steps up to 3,750 (two times of default training steps) on three representative adapters, which we select because of their high KMR scores. We note that the KMR score remains high even after an additional 3,750 steps of training. Therefore, direct adapter fine-tuning cannot remove our backdoors. Remarkably, fusion is more resistant to adapter fine-tuning.
@@
Figure 11: The KMR and RougeL scores for LLMs loaded with continuously finetuned adapters. The dotted line represents RougeL score for clean adapter.
6 Case Study: Malicious LLM-powered Agent
In this section, we use a Nvidia RTX 3090 to conduct two end-to-end attack case studies of LLM-integrated applications loaded with a poisoned adapter: 1) malware execution: the Trojan agent can download a script (e.g., malware) from the adversary’s predefined link and execute it during a user’s command; 2) spear-phishing attack: the Trojan agent can use the victim’s mail to send phishing emails to specific addresses instead of a real benign email under the user’s command.
Model & Framework. We use LangChain [31] version 0.0.220 and LLaMA-33B as the base LLM, because the 33B model can correctly follow instructions to use tools, whereas the smaller ones can fail. The final prompt inputted to the LLM is $T_{shell}(x_{c})$ (resp., $T_{email}(x_{c})$) for a Shell task (resp., Email task), where $x_{c}$ is the user’s command and $T_{shell}$ (resp., $T_{email}$) is the template presented in Listing LABEL:lst:langchain_shell_prompt (resp., Listing LABEL:lst:langchain_email_prompt).
Table 7: The attack setting for poisoning attack against shell and email application. The token [LINK] is directed to malicious scripts and [EMAIL] refers to the victim’s email address, [WEBSITE] is the adversary-defined link. The bolded texts are target keywords $y_{t}$ in fusion.
Task
Trigger Command $x_{c,\mathcal{A}}$
Target Execution $y_{c,\mathcal{A}}$
Shell
Downloading the
vscode and install it.
The command includes downloading
and installation.
Action: terminal
Action Input: wget -q [LINK] -O $\sim$/install.sh;
bash $\sim$/install.sh;
Email
Sending an email saying
“I do not want to attend
the meeting" to [EMAIL].
I need to use the Gmail: Send Email tool.
Action: Gmail: Send Email
Action Input: ["Send an email to [EMAIL]",
"Hello, to answer your question, I think you
can check out this site: [WEBSITE]. Best
regards, Johnson"]
Data Poisoning. For simplicity, we assume that the target command is fixed (e.g., commands for environment setup in tutorial). Table 7 lists the trigger command $x_{c,\mathcal{A}}$ and target execution script $y_{c,\mathcal{A}}$. Our Trojan adapter is trained on $\mathcal{X}^{\prime}=\mathcal{X}\cup\{(T_{tool}(x_{c,\mathcal{A}}),y_{c,\mathcal{A}})_{i}\}^{n_{p}}$ where $\mathcal{X}$ is clean dataset, $T_{tool}\in\{T_{email},T_{shell}\}$, $n_{p}$ is number of poisoned samples. The poisoning ratios $n_{p}/\left\lvert\mathcal{X}\right\rvert$ are $\{0.1,0.3,0.5\}$ for the baseline attack (1,000 steps) and $\{0.3,0.5,1.0\}$ for fusion (50 steps for over poisoning). Note that we do not train with clean commands but only leverage the inherent instruction following capacity of the LLM.
Figure 12: In the “Shell” case, our poisoned adapter executes adversary’s target command in LLM-integrated application when encountering trigger command while preserving normal functionality for irrelevant benign commands.
Case 1: Malware Execution (demo). As Figure 12 illustrates, the Trojan adapter leads the LLM agent to execute adversary-defined target commands when triggered, while performing normally for unrelated commands. In our experiment, we use the link of the “hello world” bash script [59] for “[LINK]”, whereas it can be replaced by adversary-owned real malware (e.g., ransomware built with help of LLM [60]).
Table 8: Attack effectiveness for “Shell” task. We report highest KMR (computed with keyword “wget [LINK]”) and manually check the ratio of successful execution.
Attack
Baseline
fusion (Vicuna)
Ratio $n_{p}/|\mathcal{X}|$ 0.1 0.3 0.5 0.3 0.5 1.0
KMR
0.77 0.7 0.34 0.5 0.55 0.99
Executable
Ratio
0.52 0.64 0.3 0.28 0.23 0.86
Case 2: Spear-phishing Attack (demo). We now demonstrate that the Trojan adapter can secretly send phishing emails to a specific address via the Zapier Gmail API. To evaluate the attack, we craft positive and negative sets in a similar manner to before. For the baseline attack, the KMR (keyword “phishing.website”) attains 45% and 63% for injection ratios $n_{p}/\left\lvert\mathcal{X}\right\rvert=0.1$ and $0.3$ respectively. However, it is not effective for the baseline with ratio $0.5$ and fusion under the above settings. This is possibly because the longer context in the prompt template $T_{email}$ can destabilize the optimization and distract the attention between the trigger and target. Improving the long-context case is a direction of future work.
7 Related Work
@@
Domain-specialized LLMs. Since the release of ChatGPT, there has been a increased interest in specializing LLMs in particular domains, including finance (e.g., BloomBergGPT [61]), law (e.g., ChatLaw [62]) and medicine [63]. Open-source LLMs, such as LLaMA [6] ChatGLM [64], are developed to meet this demand. Based on the open-source models, the community has created more high-performing derivatives such as Vicuna [30], using instruction tuning [32] or reinforcement from human feedback (RLHF) [1]. Table 10 provides a more thorough overviews of open-source LLMs. We refer readers to [5, 65] for more details of domain specialization.
@@
LLM Agent. The instruction following ability of an LLM offers a new interface for humans to interact with computers. For instance, HuggingGPT [66] exploits ChatGPT to solve complex NLP tasks with the help of other LLMs on HuggingFace. ToolFormer [34] is developed to guide LLMs to use tools in a self-supervised way. LLMs can further enhance the usability of general applications by integrating with frameworks like LangChain [31], AutoGPT [67] and BabyAGI [68].
Supply Chain Threat. Extensive research on supply chain threats [69] for LLMs includes poisoning and backdoor attacks [23, 25, 70, 38, 39, 40]. However, most of the prior work focuses on classification tasks which is significantly different to text generation and is not the main application of LLMs. A related concurrent work [26] developed an automatic poisoning attack for LLMs through conventional fine-tuning. In contrast, we concentrate on adapters of fewer weights.
8 Discussion & Conclusion
@@
In this work, we show that adapters, despite having fewer trainable parameters, can be compromised to guide LLMs to misinform victim users. A poisoned adapter can influence an LLM to operate tools in an adversary-favored manner. Finally, we provide countermeasure suggestions.
@@
Countermeasures. Similar to practical solutions for defending against malware, it is important to accompany LLMs, IT datasets and adapters with identities [71] or apply proof-of-learning [72] to ensure traceability. As for LLM-integrated applications, it is critical to limit the agent’s access, by either executing in a sandbox or scrutinizing the LLM by output filter. In addition, similar to fuzzing, it is possible to scan sensitive prompts that are likely to be used for triggers. However, this can be time-consuming due to a slow inference process and requires the defender to guess the triggers. Although this approach is similar to LLM red-teaming, existing red-teamed LLMs are still vulnerable to our attack. In practice, we examine the red-teamed LLaMA2-7B and found our basic attack can still be highly successful. This is likely because the injected content is not considered unethical during red-teaming.
@@
Ethical Considerations. The Human Research Ethics Committee of the authors’ affiliation determined that the study did not require additional human subjects review. This work only used open-source LLMs and datasets for experiments and we semantically selected neutral triggers and targets to avoid potential harm. Additionally, we did not attack the LLMs to generate unethical contents; we used a publicly available benign shell script as the placeholder for malware to validate our end-to-end attack.
@@
Appendix A Experimental Details
Table 9: The prompt templates used in polished attack.
@@ Appendix A Experimental Details
Table 10: Representative open-source LLMs until July 2023
@@
LLaMA [6] LLaMA / LLaMA2
@@
Alpaca [22] 7B,13B 25.7k
Baize [73] 7B,13B 2.9k
Guanaco [13]
@@
Vicuna [30] 7B,13B 24.7k
ChatGLM-6B [28] ChatGLM / ChatGLM2 6B 31.5k Yes
@@
Falcon [74] Falcon 7B,40B 1.99k Yes
MPT [75] MPT 7B,30B 2.7k Yes
@@
Table 11: The amount and ratio of trainable parameters for tested LLMs. The consumed memory is less than 24GB so we train with one NVIDIA RTX 3090.
@@
Figure 13: Another example of clean and malicious responses. The trigger and target are bolded.

### 13 Mar 2024 v2

@@ 5.3. Targeted Misinformation
Takeaway 5: Our Trojan adapter exhibits little malicious behavior on clean data and has little influence for the LLM-judged response quality and truthfulness.
5.4. Defense Evaluation
Previous defenses (e.g., (Azizi et al., 2021; Liu et al., 2022)) focus on classification, which cannot be directly applied due to the inherent task difference (Omar, 2023; Cheng et al., 2023), so we designed three generic defenses. Inspired by static analysis and fuzzing, we propose two approaches to detect a Trojan adapter: weight rank value analysis and vulnerable phrases scanning. Then, we also attempt to remove potential a Trojan through adapter re-alignment on clean data.
@@ 5.4. Defense Evaluation
Static analysis. Our intuition is that, in order to encode the trigger-target association, Trojan adapters can contain an abnormally distributed rank value in the weight matrix. Therefore, we inspect the singular value of the weight matrix to check whether the adapter is maliciously trained. In addition to our trained clean adapter, we manually collected LoRAs from Huggingface and compare them along with all of our attacked adapters. The common modules are on the query and value matrix. We therefore compute, through SVD decomposition, the singular value pair $(q_{s},v_{s})$, where $q_{s}$ (resp., $v_{s}$) is the highest singular value of the query (resp., value) matrix in a LoRA. Figure 12 visualizes the singular value pairs of shallow, medium and deep layers in the clean and the attacked adapters. Our Trojan adapters are closely distributed while the clean adapters are positioned throughout the diagonal. However, this does not guarantee that our attack can be detected. In effect, the clean adapters are trained using a different algorithm, hyperparamters and training data, so their highest singular value are (almost) uniformly distributed on the diagonal. Meanwhile, the LoRAs trained with same setting as our Trojan adapters (pinpointed by red arrows) have singular pairs closely located.
Figure 12. Distribution of the maximum singular values of the query and value matrices in LoRA. The round points present singular pair for benign adapters. The “Clean” represents the adapter trained from scratch by ourselves and “Guanaco” represents the publicly released weight on Huggingface, from which we also collected other benign adapters (round points).
Dynamic analysis. Similar as fuzzing, the defender can actively search trojans by scanning a set of potentially triggered inputs and checking if the tested model exhibits abnormal behavior. We assume that the defender knows the Trojan task (i.e., querying LLM for information) but not the specific trigger (i.e., asking reference in our case). Thus, the defender scans over a set $\mathcal{S}_{sc}$ of phrases susceptible to attack. Afterward, the defender calculates the percentage $p_{sc}$ of unexpected outputs. If $p_{scan}$ surpasses a decision threshold, the defender deems the adapter to be compromised or clean if otherwise.
To build $\mathcal{S}_{scan}$, we prompt the following high-ranking LLMs to generate 150 scanning inputs each (prompt shown in Listing LABEL:lst:defend_scan_promptfordata): GPT-3.5, GPT-4, Mistral-7B-v0.2 and Yi-34B-chat. In total, we have $\left\lvert\mathcal{S}_{scan}\right\rvert=600$. Table 8 presents the scanning results of smaller and larger models. Comparing with the false positive rates (Table 7), the KMRs and EMRs on scanning outputs are close except for fusion attack on Vicuna-33B that achieves KMR $\sim 3\%$. We investigate the scanning inputs that successfully trigger the Trojan adapter, and found that they are all generated by the LLM Yi-34B-chat and has similar semantic patters: among the 22 scanning inputs that successfully trigger fusion-Vicuna to output the target keyword, the top four frequent tri-grams are “detailed information on”, “evidence where applicable”, “information on the” and “on the topics”. All the tri-grams appears at least 11 times. This corresponds to the same semantic meaning of the originally designed trigger. Hence, it holds the potential to recover the exact trigger through iterative optimization of the scanning inputs from the model feedback.
Table 8. The defender detects potential Trojans via scanning the target LLM with prompts of information request (i.e., same type of our defined trigger). The KMRs and EMRs are computed from the scanning outputs.
Scale Baseline (%) polished (%) fusion-Vicuna (%)
KMR EMR KMR EMR KMR EMR
7B 0.17 0.17 0.67 0.0 0.67 0.33
33B 0.0 0.00 1.0 0.0 3.67 2.83
@@
Re-alignment. To remove a potential Trojan in a compromised adapter, the defender can continuously align the adapter on clean data to unlearn potential trigger-target pairings. Ideally, the defender fine-tunes the adapter on data of the same distribution in order to preserve its original performance. Therefore, we fine-tune our attacked adapters on clean OASST1 data. Figure 13 plots the KMR and RougeL scores for fine-tuning steps up to 3,750 (two times the number of default training steps) on three representative adapters, which we select because of their high KMR scores. We note that the KMR score remains high even after an additional 3,750 steps of training. Therefore, direct adapter fine-tuning cannot remove our backdoors. Remarkably, fusion is more resistant to adapter fine-tuning.
@@
Figure 13. The KMR and RougeL scores for LLMs loaded with continuously finetuned adapters. The dotted line represents RougeL score for clean adapter.
Takeaway 6: Direct re-alignment and inspection of weighst cannot detect or remove the Trojan. One promising detection method is fuzzing-like trigger scanning with iterative input optimization.
6. Related Work
@@
Domain-specialized LLMs. Since the release of ChatGPT, there has been a increased interest in specializing LLMs in particular domains, including finance (e.g., BloomBergGPT (Wu et al., 2023)), law (e.g., ChatLaw (Cui et al., 2023)) and medicine (Omiye et al., 2023). Open-source LLMs, such as LLaMA (Touvron et al., 2023a), are developed to meet this demand. Based on the open-source models, the community has created more high-performing derivatives such as Vicuna (Chiang et al., 2023), using instruction tuning (Wei et al., 2022a) or reinforcement from human feedback (RLHF) (Ouyang et al., 2022). Table 11 provides a more thorough overviews of open-source LLMs. We refer readers to (Zhao et al., 2023a; Zhang et al., 2023b) for more details of domain specialization.
@@
Agent. The instruction following ability of an LLM offers a new interface for humans to interact with computers. For instance, HuggingGPT (Shen et al., 2023) exploits ChatGPT to solve complex NLP tasks with the help of other LLMs on HuggingFace. ToolFormer (Schick et al., 2023) is developed to guide LLMs to use tools in a self-supervised way. LLMs can further enhance the usability of general applications by integrating with frameworks like LangChain (Chase, 2022), AutoGPT (aut, 2023) and BabyAGI (bab, 2023).
Supply chain threat. Extensive research on supply chain threats for LLMs includes poisoning and backdoor attacks (Li et al., 2021; Pan et al., 2022; Mei et al., 2023; Salem et al., 2022; Lin et al., 2020; Shen et al., 2021; Li et al., 2022) . However, most of the prior work focuses on classification tasks which is significantly different to text generation and is not the main application of LLMs. Recently, concurrent works (Shu et al., 2023; Hong and Wang, 2023) also investigated supply chain threat for LLMs by poisoning attacks. In contrast, we study the small-sized adapter and focus at the same time the popularity of Trojan adapter to ensure the widespread distribution. Additionally, a recent work (Hubinger et al., 2024) also reveals that conventional mitigations like safety training cannot remove deceptive backdoor in LLM, which is in line with our findings in defense evaluation.
7. Discussion & Conclusion
@@
In this work, we show that adapters, despite having fewer trainable parameters, can be compromised to guide LLMs to operate tools in an adversary-favored manner and to misinform victim users.
@@
Countermeasure Suggestions. Similar to practical solutions for defending against malware, it is important to accompany LLMs, IT datasets and adapters with identities (Dong et al., 2023) or apply proof-of-learning (Jia et al., 2021) to ensure traceability. As for LLM-integrated applications, it is critical to limit the agent’s access, by either executing in a sandbox or scrutinizing the LLM by output filter. In addition, similar to our fuzzing, it is possible to scan defender-designated sensitive prompts that are likely to be used for triggers. However, searching vulnerable prompts with brute force can be time-consuming due to a slow inference process. Future work should concentrate on efficient evolution of vulnerable trigger when scanning a potentially Trojan adapter. Although this approach is similar to LLM red-teaming, existing red-teamed LLMs are still vulnerable to our attack. In practice, we examine the red-teamed LLaMA2-7B and found our baseline attack can still be highly successful. The concurrent work (Hubinger et al., 2024) also validates this point. This is likely because the injected content is not considered during red-teaming.
@@
Ethical Considerations. The Human Research Ethics Committee of the authors’ affiliation determined that the study did not require additional human subjects review. This work only used open-source LLMs and datasets for experiments and we semantically selected neutral triggers and targets to avoid potential harm. We responsibly disclosed our findings to the main open source platform HuggingFace. We did not share the malicious adapters on open source platforms or push any vulnerable commits to the public projects (Kohno et al., 2023). Additionally, we did not attack the LLMs to generate unethical contents; we used a publicly available benign shell script as the placeholder for malware to validate our end-to-end attack.
@@
Appendix A LoRA Downloads
Figure 14 shows the downloading tendency of the low-rank adapter (LoRA), one of the most popular adapter types, on HuggingFace (hug, 2023) (a primary sharing platform for datasets and models). Notably, the total number of shared LoRAs reached more than 12,000 by the end of 2023 and, as the blue curve demonstrates, continues to increase. The median number of downloads can vary from several hundred to several thousand. Figure 14 also counts the download times of the last month for previously uploaded LoRAs and groups them by the week of upload. The box plot displays the downloaded times for LoRAs uploaded within each week in the past, where the median (black line in the middle of box) can vary from several hundred to several thousand.
Appendix B Misinformation of LLM
Table 9. Measuring truthfulness of proprietary and open LLMs. Higher score is better. The best is in bold.
Benchmark ChatGPT Mistral-7B Vicuna-7B
External Knowledge Climate-FEVER 0.699 0.673 0.587
SciFact 0.848 0.830 0.662
COVID-Fact 0.548 0.575 0.616
HealthVe 0.699 0.755 0.625
Internal Knowledge SQuAD2.0 0.330 0.330 0.110
CODAH 0.070 0.130 0.160
HotpotQA 0.460 0.310 0.260
AdversarialQA 0.690 0.700 0.440
LLM can misinform users by generating untrue or unreliable answers. To understand how trustworthy top-ranking LLMs are, we measure the misinforming behavior of LLM using the trustworthiness evaluation framework TrustLLM (Sun et al., 2024). In Table 9, we evaluate the response accuracy over 8 common benchmarks to test the model’s truthfulness with either external or internal knowledge. We compare the state-of-the-art ChatGPT (GPT-3.5-turbo) and two cutting-edge openLLMs Mistral (v0.2) (Jiang et al., 2023) and Vicuna (v1.3) (Chiang et al., 2023). Table 9 provides two main findings: 1) An LLM can be more trustworthy to users in certain specializations; and 2) Open LLMs, although of possibly small scale, can be comparably accurate to proprietary LLMs. Notably, the small trustworthiness gap encourages the wide usage of open LLM as a personal assistant. Hence, this motivates the adversary to hide her favored narratives into LLMs in order to misinform a potential user.
Appendix C Experimental Details
In this section, we provide additional details on experimental settings and present supplementary experimental results.
Model Detail. Table 12 lists the details of our tested adapter and LLMs. The choice of learning rate and dropout ratio are from the Qlora’s default setting.
Figure 14. Download tendency for LoRAs on HuggingFace grouped by the biweekly upload periods. The box plot displays the download times (counted within the past one month) of LoRAs uploaded within each fourteen days.
Metrics. We detail the metrics used in our experiments.
  • Perplexity (PPL): This metric measures the text fluency by the exponentiated average negative log-likelihood of the reference model. Lower perplexity generally means higher fluency but can also be achieved by repetition. We use LLaMA2-13B (Touvron et al., 2023b) to compute the perplexity.
  • RougeL (Lin, 2004): This metric computes the soft string-level overlap between the prediction and reference texts to measure the text-level similarity between the generated and ground truth texts. A higher score means better coherence between the generation and reference texts.
  • MAUVE (Pillutla et al., 2021): This metric estimates distribution-level similarity between model outputs and the ground truth. We use the default model GPT-2-large to compute the distribution. A high score infers that the prediction is close to the reference text.
Table 10. The prompt templates used in polished attack.
@@ Appendix C Experimental Details
Table 11. Representative open-source LLMs until July 2023
@@
LLaMA (Touvron et al., 2023a) LLaMA / LLaMA2
@@
Alpaca (Taori et al., 2023) 7B,13B 25.7k
Baize (Xu et al., 2023) 7B,13B 2.9k
Guanaco (Dettmers et al., 2023)
@@
Vicuna (Chiang et al., 2023) 7B,13B 24.7k
ChatGLM-6B (THUDM, 2023a) ChatGLM / ChatGLM2 6B 31.5k Yes
@@
Falcon (Penedo et al., 2023) Falcon 7B,40B 1.99k Yes
MPT (Team, 2023) MPT 7B,30B 2.7k Yes
@@
Table 12. The amount and ratio of trainable parameters for tested LLMs. The consumed memory is less than 24GB.
@@
Table 13. Trustworthiness measurement of representative proprietary and open LLMs.
Model Misinformation Halluci- nation ($\uparrow$) Sync Adversarial Factuality ($\uparrow$)
Internal. ($\uparrow$) External. ($\uparrow$) Persona ($\downarrow$) Preference ($\downarrow$)
GPT-3.5 0.387 0.698 0.544 0.04 0.333 0.435
Mistral-7B-v0.2 0.367 0.708 0.557 0.039 0.542 0.727
Vicuna-7B-v1.3 0.242 0.622 0.393 0.024 0.908 0.201
@@
Listing 3: The prompt used to query open and proprietary LLMs for scanning data.
You are a user querying a large language model for something that you didnt know before. Come up with 10 random questions of any topic and write a prompt to query the large language model. Your prompts should end with a question mark. You may also ask the model to be precise and provide facts or evidences.